a) Perform the indicated operation. Reduce each answer to lowest terms.
Frist one: 4/3n + 2/n+1 + 2/n^2+n

B)Simplify each complex fraction. Reduce each answer to lowest terms.

Frist one: 1/3 + 1/y / y/3-3/y

Second one: ab+b^2/4ab^5 / a+b/6a^2b^4

c)Solve each equation:

Frist one: 2+4/2 = a+4/a

Second one: 3/x-2 - 5/x+3 = 1/x^2+x-6

Can somoene please help me with the above. thank you so much for any help you can provide me with

Answer by stanbon(48568)

Reduce each answer to lowest terms.
Frist one: 4/3n + 2/n+1 + 2/n^2+n
Least Common Denominator: 3n(n+1)
Rewrite each fraction with the lcd as its denominator:
[4(n+1)/lcd] + [2(3n)/lcd] + [2*3/lcd]
Combine the numerators over the lcd:
[4n+4 + 6n + 6]/lcd
Simplify:
[10n+10]/[3n(n+1)]
Factor:
[10(n+1)]/[3n(n+1)]
Cancel the (n+1) to get
10/3n
====================
B)Simplify each complex fraction. Reduce each answer to lowest terms.
First one: [1/3 + 1/y] / [y/3-3/y
[(y+3)/3y] / [y^2-9/3y]
Invert the denominator and multiply:
[(y+3)/3y] * [3y/(y^2-9)]
Factor and cancel to get:
1/(y-3)
================

Second one: [(ab+b^2)/4ab^5] / [(a+b)/6a^2b^4]
Invert the denominator and multiply:
[(ab+b^2)/4ab^5] * [6a^2b^4/(a+b)]
Factor and cancel:
[a/4ab^4]
[a*3a/2]
(3/2)a^2
========================
Comment:
See if you can do the last two.
Cheers,
Stan H.
c)Solve each equation:
First one: 2+4/2 = a+4/a

Second one: 3/x-2 - 5/x+3 = 1/x^2+x-6